# UAV See, UGV Do: Aerial Imagery and Virtual Teach Enabling Zero-Shot Ground Vehicle Repeat

### v1

@@ V Discussion and Lessons Learned
Enabling autonomous control via the T&R framework using NeRF-derived submaps as the sole input proved challenging in multiple respects. A significant observation was the difficulty in generating point clouds and meshes with minimal noise and accurate geometry under challenging snowy conditions. NeRF model training noticeably suffered in environments where large white swaths of untouched snow resulted in minimal detail and reduced depth information in the scene. Earlier aerial data captures of the same environments produced much more detailed NeRF models when snow was not present. Interestingly, although VirT&R was capable of running without issue in all the environments considered in this paper, attempts to use VirT&R on the Grassy Loop—another route commonly used for T&R testing [5]—resulted in rapid and unpredictable localization drift in the snowy, highly unstructured environment. This highlights a potential pitfall of the method and will be further investigated through the substitution of NeRF with other volumetric rendering techniques. Additionally, the methodology used to evaluate the performance of VirT&R relative to the virtual teach path derived from the NeRF-generated mesh is subject to inherent errors, both from how the Warthog is paused to take measurements and from positioning the jig to measure the marks on the ground. It is likely that if these errors can be reduced, the evaluated performance would more closely approach that of LT&R. There remains room for improvement in the VirT&R evaluation methodology, potentially by using the geo-registered NeRF mesh to obtain pseudo-GPS coordinates for the virtual path.
@@ VI Conclusion
In this work, we presented Virtual Teach and Repeat: an autonomous route following stack capable of driving in various environments using only monocular commercial drone imagery and a piloted path in simulation as the only inputs to the system. The NeRFs of captured scenes allowed for rich point clouds and visually accurate meshes to be extracted and used for localization and pilot simulation. This extension to Teach and Repeat offers the advantages of constructing a large environment once—without requiring the physical presence of a human or UGV—and subsequently piloting any desired route within that environment without needing to capture it again. We compared the path-tracking performance of NeRF-derived submaps to LiDAR on four different routes and completed over 12 km of uninterrupted autonomous driving. VirT&R performed well in the tested environments, achieving measured RMSE values of 19.5 cm and 18.4 cm and maximum errors of 39.4 cm and 47.6 cm. Relative repeat path analysis corroborates the qualitative observation that VirT&R repeats its path smoothly, despite moving cars, pedestrians, and slight environmental changes. Finally, we intend to improve the methodology used for assessing VirT&R’s accuracy and to ensure its functionality remains robust in increasingly unstructured environments by investigating alternative methods for volumetrically rendering scenes from aerial imagery. This will further enhance Virtual Teach and Repeat, expanding cross-modal sensor support and extending the capabilities of the Teach and Repeat framework.

### v2

@@ V Discussion and Lessons Learned
Enabling autonomous control via the T&R framework using virtually-generated submaps proved challenging in multiple respects. A significant observation was the difficulty in scene reconstruction with minimal noise and accurate geometry in challenging snowy conditions. NeRF model training noticeably suffered in environments where large swaths of untouched snow resulted in minimal detail and reduced depth information. Earlier aerial surveys of the same scenes produced much more detailed NeRF models when snow was not present. Interestingly, although VirT&R was capable of running without issue in all the environments considered in this paper, attempts to use VirT&R on the Grassy Loop—another route commonly used for T&R testing [5]—resulted in unpredictable localization in the flat, highly unstructured environment. This highlights a potential pitfall of the method and will be investigated further through the substitution of NeRF with other scene reconstruction techniques, such as classic photogrammetry. Additionally, the methodology used to evaluate the performance of VirT&R is subject to inherent errors, both in how the UGV is paused to take measurements and in the positioning of the jig to measure the offsets. It is likely that if these errors could be reduced, the evaluated performance would more closely approach that of LT&R. There remains room for improvement in the VirT&R evaluation methodology, as the current experimental structure does not lend itself to being evaluated conventionally with GPS. To reduce the compounded errors in the ‘pseudo-GPS’, fiducial tags, existing 3D models like Google Earth, or ground control points with known absolute GPS information could be used to further anchor the reconstruction in the same frame as GPS data taken during live VirT&R repeats.
@@ VI Conclusion
In this work, we presented Virtual Teach and Repeat: an autonomous route following stack capable of driving in various environments using only monocular commercial drone imagery and a piloted path in simulation as the only inputs to the system. The NeRFs of captured scenes allowed for rich point clouds and visually accurate meshes to be extracted and used for localization and pilot simulation. This extension to Teach and Repeat offers the advantages of constructing a large environment once—without requiring the physical presence of a human or UGV—and subsequently piloting any desired route within that environment without needing to capture it again. We compared the path-tracking performance of NeRF-derived submaps to LiDAR on four different routes and completed over 12 km of uninterrupted autonomous driving. VirT&R performed well in the tested environments, achieving measured RMSE values of 19.5 cm and 18.4 cm and maximum errors of 39.4 cm and 47.6 cm. Relative repeat path analysis corroborates the qualitative observation that VirT&R repeats its path smoothly, despite moving cars, pedestrians, and slight environmental changes. Finally, we intend to improve the methodology for assessing VirT&R’s accuracy and to test in increasingly unstructured environments. This will further enhance Virtual Teach and Repeat, expand cross-modal sensor support, and extend the capabilities of the Teach and Repeat framework.